OpenStudy (anonymous):

One number is 1 less than a second number. Twice the second number is 19 less than 5 times the first. Find the two numbers.

OpenStudy (anonymous):

Sorry, I should have clarified. I'm working on solving systems of linear equations by the substitution method.

OpenStudy (nurali):

One number is 1 less than a second number" ___ f = s - 1 "Twice the second number is 19 less than 5 times the first" ___ 2s = 5f - 19 substituting the value of f ___ 2s = 5(s - 1) - 19 2s = 5s - 5 - 19 24 = 3s 24/3 = s s=8 f = (8) - 1 = 7